OpenStudy (anonymous):

how do you do the steps to do long division of polynomials for this equation??(x^2+x-3) / (x-2)

OpenStudy (anonymous):

There are two methods: 1. factor the numerator and then reduce the fraction; 2. use the following property: (a + b + c)/d = a/d + b/d + c/d. There is also a third method called synthetic division that can be used only when the divisor polynomial is written in the form x - r, where "r" is any real number. In your case, is it much easier to apply the third method as you divisor is in the form required, x - 2.

OpenStudy (anonymous):

There are two methods: 1. factor the numerator and then reduce the fraction; 2. use the following property: (a + b + c)/d = a/d + b/d + c/d. There is also a third method called synthetic division that can be used only when the divisor polynomial is written in the form x - r, where "r" is any real number. In your case, is it much easier to apply the third method as you divisor is in the form required, x - 2.